Use Kepler’s first law to derive explicit expressions for the perihelion and aphelion distances for a planet in terms of the semi-major axis a, and the eccentricity e, for an elliptical orbit. r=a(1-e^2)/1+ecos(theta) Calculate these distances for Mercury. What are the perihelion and aphelion speeds of Mercury? V^2=GM(2/r-1/a)
Answer #1

Solution attached.Solution the Orbital period of fleecury is 87.97 days. the planets Semi major anu à ai pe=a3 the gemi masor anie in a(AU) =Closet at perikelion, 0-0 r =all-e) = 1 AU*(1-0.017). = 0.983 AU IC te) = 1 AU*(1+0.017) Sant = 1.017 AU va - amrage speed aThe total eneegy, E = 1 mv-amom E= - Grom za a encegy values divide to get the vz-AMO --CM for speed. Some v- samole valoda)